HIGH-RESOLUTION INFRARED SPECTROSCOPY OF HCl DIMER IN SOLID PARAHYDROGEN

\underline{DAVID T. ANDERSON}, Department of Chemistry, University of Wyoming, Laramie, WY 82071-3838; SIMON TAM, and MARIO E. FAJARDO, Propulsion Directorate, US Air Force Research Laboratory, AFRL/PRSP, Edwards Air Force Base, CA 93524-7680.

The infrared spectroscopy and dynamics of HCl hydrogen bonded clusters in solid parahydrogen were investigated using high-resolution FTIR spectroscopic methods. This paper will focus on the HCl dimer species. Transitions have been observed and assigned to both the “donor” and “acceptor” HCl stretching modes of (HCl)_{2}. The spectroscopy indicates that end-over-end rotation of the dimer is quenched in solid parahydrogen, but that low frequency motions involving K rotation and possibly tunneling are still occurring. More detailed analysis is underway and will be presented at the meeting.